# Minimalist Market Design: A Framework for Economists with Policy Aspirations

### v2

@@ 7.9.2 The Global Rise of Kidney Exchange Programs Since 2008
Due to the limited size of their national pools, Austria and the Czech Republic—both of which introduced national kidney exchange systems in 2011—consolidated their programs in 2015. This collaboration led to the first cross-border kidney exchange in 2016 (Böhmig et al., 2017) and was further expanded to include Israeli centers in 2019. Between 2011 and 2019, these systems led to 81 transplants, including 33 after 2015 under the unified program, of which 8 were via cross-border exchanges (Viklicky et al., 2020).
Similarly, Italy, Spain, and Portugal established their national kidney exchange programs in 2006, 2009, and 2010, respectively. In 2017, the three countries formalized a cooperation protocol, resulting in their first successful cross-border exchange in 2018 (Valentín et al., 2019). Unlike other multi-country collaborations, however, each nation continues to operate its own independent national system before pursuing cross-border matches.
Since 2008, collaborations among the European market design community, clinicians, and policymakers have driven significant advances in kidney exchange under the guidance of market design experts such as Tommy Andersson, Peter Biró, David Manlove, and Ana Viana. A major milestone was the establishment of the European Network for Collaboration on Kidney Exchange Programs (ENCKEP), supported by the European Cooperation in Science and Technology (COST) from September 2016 to March 2021 (Biro et al., 2019; ENCKEP, 2024). Among ENCKEP’s achievements is the development of KEP-SOFT, a dedicated software platform facilitating transnational kidney exchange efforts in Europe (KEP-SOFT, 2024).
Further affirming the growing enthusiasm for kidney exchange in Europe, the European Commission launched a 2023 call for proposals for action grants under its Health Systems and Healthcare Workforce initiative, specifically targeting “Facilitating Organ Paired Exchange” (European Commission, 2023). This initiative seeks to harmonize organ donation and transplantation practices across EU member states by developing a shared algorithm for matching donor-patient pairs and establishing common protocols, platforms, and governance frameworks. The ultimate goal is to create an EU-wide system that expands kidney exchange opportunities and improves patient outcomes.
Another significant milestone came in 2024 with the advocacy of market design experts Itai Ashlagi, Ágnes Cseh, David Manlove, and Axel Ockenfels. Germany, previously one of the last European countries where kidney exchange was illegal, legalized the practice through an amendment to its Transplantation Law (Ashlagi et al., 2024; Ockenfels et al., 2024).
@@ 8.3 Switzer’s Reform Proposal for the USMA-2006 Mechanism
Under Switzer’s proposed priority structure, all units within a given branch use the same criteria for allocation, with price always being the primary consideration. This specific design choice was the source of my skepticism regarding the mechanism’s viability for the Army. As discussed in Section 8.2, the Army’s policy maintained a delicate balance by granting BRADSO volunteers elevated priority for only a fraction of available positions under the USMA-2006 mechanism, not all. At the time, although I didn’t know the exact rationale behind this limitation, I strongly believed it wasn’t arbitrary. [^90] As evidenced by Footnote 88, this hunch turned out to be correct. This is why I viewed this aspect of Switzer’s mechanism as a potential “deal-breaker” and chose to refine it by incorporating both criteria, as the Army had done. In my view, while Switzer’s mechanism successfully addressed some critical shortcomings of the USMA-2006 mechanism, it would likely introduce new ones.
@@ 9 Design of Property Rights
Despite the theoretical interest in choice rules, our use-inspired research had never interfered with institutional choice rules—unless they were among the root causes of failures in the underlying mechanisms—up until the early 2010s. The minimalist mindset that guided our holistic approach to both research and policy was a key reason for this stance. This remained true until Boston Mayor Thomas Menino’s pivotal 2012 State of the City Address, where he pledged to increase the fraction of students assigned to schools in their neighborhoods. [^96] One instance where a choice rule directly contributed to failures was the ROTC branching mechanism, leading to a proposed mechanism with an alternative choice rule, as discussed in Section 8.6. From a pragmatic standpoint, interfering with choice rules would not only deviate from minimalist market design but also risk alienating interest groups negatively impacted by our proposals.
@@ 9.1 Subtleties of Reserve Systems
When reserve systems were first studied by Hafalir et al. (2013) in the market design literature, they presented them directly as the minimum guarantee choice rule, [^97] Hafalir et al. (2013) refer to the minimum guarantee choice rule policy as soft reserves. without discussing other possible variants. The significance of the processing sequence of positions—the order of precedence—was first introduced and studied by Kominers and Sönmez (2013, 2016) within a more general framework that incorporates slot-specific priorities into the matching with contracts model discussed in Section 8.4. Building on this use-inspired basic theory, which became pivotal in several subsequent “discovery-invention cycles,” Dur et al. (2018b) examine our current real-world setting as a special case of this model.
@@
This subtlety in the order of precedence in reserve systems—and particularly the distinction between the over-and-above and minimum guarantee policies—is so elusive that it has led to unintended consequences in several real-world applications. In some cases, decision makers have completely overlooked this distinction, leaving major loopholes in systems that otherwise meticulously control parameters such as the fraction of reserved units. [^98] See the The Hindu op-ed Pathak and Sönmez (2019) for an example of such a loophole when the 103rd Constitutional Amendment Act, 2019 introduced reservations for economically weaker sections of society. Occasionally, politically motivated actors have exploited these loopholes through reforms with significant distributional consequences. [^99] See Pathak et al. (2020a) for an example of this, when officials of the U.S. Department of Homeland Security reversed the processing sequence of two categories of H-1B visas in response to President Trump’s Buy American and Hire American Executive Order in 2017. In other instances, as illustrated in our next case study, system operators intended to implement either the over-and-above or minimum guarantee policy but ended up implementing the wrong one—either because they were unaware of the distinction or did not know how to operationalize the intended policy, especially in a multi-institution setting.
@@ 9.2 The Demise of Neighborhood Priority in Boston School Choice
Building on Dur et al. (2018b), our next application examines Boston’s neighborhood priority policy and the unintended consequences of its flawed implementation between 2005 and 2013, which ultimately disadvantaged applicants seeking to attend their neighborhood schools.100100100The period from 2005 to 2013 provides a conservative estimate of the duration of the flawed implementation of Boston’s neighborhood priorities. The neighborhood priority policy intended for use under the DA mechanism during this period was originally adopted in 2000 under Boston mechanism. The flawed implementation of this policy, discussed in Section 9.2.2, led to the de facto elimination of neighborhood priorities under the DA mechanism and would have caused similar unintended consequences in the Boston mechanism had an analogous error occurred. However, it remains unclear whether such a flawed implementation of the neighborhood priority policy existed between 2000 and 2005 under the Boston mechanism.
@@ 9.2.3 Phasing Out Neighborhood Priority in Boston: 2013 Reform for Transparency
Though I didn’t know it at the time, one of the regular participants in the public meetings was Peng Shi, an MIT graduate student in operations research who was auditing my market design course at Boston College. Recognizing the gridlock as an opportunity, Shi proposed an entirely new approach during one of the meetings: the “Home-Based” system (Shi, 2015).101101101Later, I learned that Shi had been studying the crisis with an eye toward helping guide a reform as early as Spring 2012. Before the Home-Based system, he also proposed other plans when the EAC elicited suggestions from the wider community in late summer. This system provided each student with a customized list of schools based on their home address, reducing the number of schools they could rank while ensuring that highly-rated options were included.102102102See Shi (2022) for an in-depth exploration of Shi’s broader methodology, which optimizes parameters such as student priority distribution within mechanisms like DA and TTC, parameters that are typically treated as exogenous in design economics. His proposal gained traction, as EAC members were eager to bridge the divide between the committee’s opposing camps.
@@ 10.1 India’s Reservation System and the Quest for Social Justice
Affirmative action for various disadvantaged groups is embedded in the 1950 Constitution of India through articles 14-18.103103103This segment of the Constitution is referred to as the Equality Code in India. In practice, these protective policies are implemented through a reservation system that sets aside a certain fraction of positions in public jobs and seats at public schools for various protected groups. Naturally, the Constitution itself does not provide any rigorous formula to implement these policies. Nevertheless, in a landmark Supreme Court judgment Indra Sawhney vs. Union of India (1992), the justices formulated these provisions in a rigorous way, at least for the more basic versions of the reservation system.104104104Considered the main reference on the reservation system in India, this case is popularly known as the Mandal Commission case.
@@ 10.2 Vertical Reservations, Migration and the Over-and-Above Choice Rule
From a technical point of view, the social identity-based vertical reservation system meant that no individual could be a member of multiple VR-protected groups. While VR protections are also provided to individuals in economic disability since 2019 with the enactment of the 103rd Amendment,105105105This more recent vertical category is called Economically Weak Segments (EWS.) members of the earlier and social identity-based VR-protected groups are excluded from its scope,106106106This exclusion is highly controversial in India. See Sönmez and Ünver (2022) for a case study in minimalist market design that analyzes this exclusion in the 103rd Amendment. thereby still maintaining the non-overlapping structure of VR-protected groups in the country.
@@
A subtle but important aspect of the above mandate is its reliance on a basic technical tool commonly referred to as migration in Indian legal terminology.107107107The concept of migration is also known as mobility in Indian legal terminology.
@@ 10.3 Horizontal Reservations, Adjustments and the Minimum Guarantee Choice Rule
This passage leaves unclear what exactly the “necessary adjustments” refer to and how they are to be made. The ambiguity surrounding these adjustments and the implementation of the HR policy was later clarified by the Supreme Court in Anil Kumar Gupta vs. State of U.P. (1995).108108108The formulation of the HR policy in Anil Kumar Gupta vs. State of U.P. (1995) is further clarified in a subsequent Supreme Court judgment, Rajesh Kumar Daria vs. Rajasthan Public Service Commission (2007). In this judgment, the Supreme Court further elaborated two key characteristics of the HR policy that distinguish it from the VR policy:
@@
Possibly to avoid complications arising from overlaps between “property rights” gained through these two types of affirmative action policies, the justices also recommended implementing the HR policy within each vertical category, including the open category.109109109This mode of HR policy is referred to as compartmentalized horizontal reservations in Anil Kumar Gupta vs. State of U.P. (1995). For example, if there are horizontal reservations for women, they are provided with preferential access for a fraction of the positions in each category (i.e., open, SC, ST, OBC, etc.) separately.
@@
Their procedure begins by disregarding the HR policy in the first two steps and focusing solely on the VR policy. In the first step, open positions are tentatively allocated using SSD, followed by the allocation of VR-protected positions to the remaining eligible individuals, again using SSD, in the second step.110110110Note that these two steps together form the over-and-above choice rule.
@@
Except for the recent EWS category, a member of a VR-protected category is not required to declare membership in their category.111111111This contrasts with the documentation needed to prove ineligibility for social identity (i.e., caste or tribal) based VR-protected categories, which is a prerequisite for claiming membership in the EWS category. This typically involves documentation of forward caste membership in India. By claiming their VR protections, an individual—unless a meritorious reserved candidate—loses their open category HR protections under the SCI-AKG choice rule. Thus, this mechanism violates a form of the incentive compatibility axiom, first formulated by Aygün and Bó (2021) in the context of affirmative action at Brazilian colleges.
@@ 10.5 External Validity for Minimalist Market Design: Life Imitates Science in a Landmark Judgment
Before its publication, the initial draft of Sönmez and Yenmez (2022) was circulated in March 2019 as a Boston College working paper Sönmez and Yenmez (2019). In this work, we established that the Supreme Court-mandated SCI-AKG choice rule had been consistently undermining intended affirmative action measures in real-world practices by failing to satisfy the no justified envy axiom.112112112In Sönmez and Yenmez (2019), the SCI-AKG choice rule is referred to as the SCI-VHR choice rule, and the no justified envy axiom is referred to as elimination of justified envy. We also documented how this failure sparked numerous litigations between 1995 and 2020, causing recurring disruptions in the country’s recruitment processes.
@@
Motivated by this case and several others that contradicted the ruling in Anil Kumar Gupta vs. State of U.P. (1995), we proposed the 2SMG choice rule in Sönmez and Yenmez (2019) as a minimalist reform of the SCI-AKG choice rule.113113113As documented in Sönmez and Yenmez (2019, 2022), the challenges arising from the failures of the Supreme Court-mandated SCI-AKG choice rule extended far beyond conflicting with high court judgments that upheld choice rules satisfying the no justified envy axiom. In many instances, state institutions using the SCI-AKG choice rule were challenged in court by VR-protected candidates due to cases of justified envy. Some of these cases were ruled in favor of the state, while others favored the plaintiffs. Additionally, there were situations where VR-protected candidates were denied their open-category HR protections even though they did not claim their VR protections. Court decisions on these challenges also varied, sometimes favoring the candidates and other times not. Moreover, in other cases, VR-protected candidates were denied open-category HR-protected positions even when no positions were reserved for their VR-protected categories.
@@
    Awarding $i$ an open-category position instead, with or without replacing another individual, does not increase the number of HR-protected positions honored within the open category.114114114Prior to Saurav Yadav vs. State of U.P. (2020), unlike the first two conditions, the necessity of this third condition as a prerequisite for the assignment of VR-protected positions remained unclear.
@@
Therefore, while the justices have not explicitly mandated the 2SMG choice rule in Saurav Yadav vs. State of U.P. (2020), they have, in effect, mandated it in the country, as no other choice rule satisfies the explicit mandates of the judgment.115115115The 2SMG choice rule is mandated in the state of Gujarat under the High Court judgment Tamannaben Ashokbhai Desai vs. Shital Amrutlal Nishar (2020).
Aligned with the version of the problem reviewed in Saurav Yadav vs. State of U.P. (2020), Theorem 12 pertains to the basic version of the problem, where all positions are identical and no individual belongs to more than one HR-protected group.116116116Strictly speaking, Saurav Yadav vs. State of U.P. (2020) does not rule out the possibility of individuals belonging to multiple HR-protected groups; it simply does not address this scenario.
@@
10.6 Overlapping HR Protections
While Persons with Disabilities (PwD) is the only HR-protected group recognized nationwide in India by a Supreme Court judgment,117117117HR protections for Persons with Disabilities were established in the Supreme Court judgment Union of India vs. National Federation of the Blind (2013). other groups, such as Women (W), receive HR protections in several states through mandates issued by High Courts or state governments. As a result, unlike VR-protected groups, individuals can often belong to multiple HR-protected groups in practical applications.

### v3

@@ 7.9.2 The Global Rise of Kidney Exchange Programs Since 2008
Due to the limited size of their national pools, Austria and the Czech Republic—both of which introduced national kidney exchange systems in 2011—merged their programs in 2015. This collaboration led to the first cross-border kidney exchange in 2016 (Böhmig et al., 2017) and was further expanded to include Israeli centers in 2019. Between 2011 and 2019, these systems led to 81 transplants, including 33 after 2015 under the unified program, of which 8 were via cross-border exchanges (Viklicky et al., 2020).
Similarly, Italy, Spain, and Portugal established their national kidney exchange programs in 2006, 2009, and 2010, respectively. In 2017, the three countries formalized a cooperation protocol resulting in their first successful cross-border exchange in 2018 (Valentín et al., 2019). Unlike other multi-country collaborations, however, each country continues to operate its own independent national system before pursuing cross-border matches.
Since 2008, collaborations among the European market design community, clinicians, and policymakers have driven significant advances in kidney exchange under the guidance of market design experts such as Tommy Andersson, Peter Biró, David Manlove, Joris van de Klundert, and Ana Viana. A major milestone was the establishment of the European Network for Collaboration on Kidney Exchange Programs (ENCKEP), supported by the European Cooperation in Science and Technology (COST) from September 2016 to March 2021 (Biro et al., 2019; ENCKEP, 2024).
ENCKEP led to a follow-on project called Software for Transnational Kidney Exchange Programmes (KEP-SOFT), funded by a COST Innovators Grant from November 2021 to October 2022. This effort produced KEPsoft, a software solution for national and international kidney exchange (KEP-SOFT, 2024). [^88] KEPsoft was jointly developed by the University of Glasgow, INESC TEC (Porto), HUN-REN-KRTK (Budapest), and Óbuda University (Budapest) (KEP-SOFT, 2024).
Further affirming the growing enthusiasm for kidney exchange in Europe, the European Commission launched a 2023 call for proposals for action grants under its Health Systems and Healthcare Workforce initiative, specifically targeting “Facilitating Organ Paired Exchange” (European Commission, 2023). This initiative aims to harmonize organ donation and transplantation practices across EU member states by developing a shared algorithm for matching donor-patient pairs and establishing common protocols, platforms, and governance frameworks. The ultimate goal is to create an EU-wide system that expands kidney exchange opportunities and improves patient outcomes. [^89] The EURO-KEP project, led by the Spanish National Transplant Organization (ONT), was the sole project funded under this call and is supported by KEPsoft software (KEP-SOFT, 2024).
@@ 8.3 Switzer’s Reform Proposal for the USMA-2006 Mechanism
Under Switzer’s proposed priority structure, all units within a given branch use the same criteria for allocation, with price always being the primary consideration. This specific design choice was the source of my skepticism regarding the mechanism’s viability for the Army. As discussed in Section 8.2, the Army’s policy maintained a delicate balance by granting BRADSO volunteers elevated priority for only a fraction of available positions under the USMA-2006 mechanism, not all. At the time, although I didn’t know the exact rationale behind this limitation, I strongly believed it wasn’t arbitrary. [^94] As evidenced by Footnote 92, this hunch turned out to be correct. This is why I viewed this aspect of Switzer’s mechanism as a potential “deal-breaker” and chose to refine it by incorporating both criteria, as the Army had done. In my view, while Switzer’s mechanism successfully addressed some critical shortcomings of the USMA-2006 mechanism, it would likely introduce new ones.
@@ 9 Design of Property Rights
Despite the theoretical interest in choice rules, our use-inspired research had never interfered with institutional choice rules—unless they were among the root causes of failures in the underlying mechanisms—up until the early 2010s. The minimalist mindset that guided our holistic approach to both research and policy was a key reason for this stance. This remained true until Boston Mayor Thomas Menino’s pivotal 2012 State of the City Address, where he pledged to increase the fraction of students assigned to schools in their neighborhoods.100100100One instance where a choice rule directly contributed to failures was the ROTC branching mechanism, leading to a proposed mechanism with an alternative choice rule, as discussed in Section 8.6. From a pragmatic standpoint, interfering with choice rules would not only deviate from minimalist market design but also risk alienating interest groups negatively impacted by our proposals.
@@ 9.1 Subtleties of Reserve Systems
When reserve systems were first studied by Hafalir et al. (2013) in the market design literature, they presented them directly as the minimum guarantee choice rule,101101101Hafalir et al. (2013) refer to the minimum guarantee choice rule policy as soft reserves. without discussing other possible variants. The significance of the processing sequence of positions—the order of precedence—was first introduced and studied by Kominers and Sönmez (2013, 2016) within a more general framework that incorporates slot-specific priorities into the matching with contracts model discussed in Section 8.4. Building on this use-inspired basic theory, which became pivotal in several subsequent “discovery-invention cycles,” Dur et al. (2018b) examine our current real-world setting as a special case of this model.
@@
This subtlety in the order of precedence in reserve systems—and particularly the distinction between the over-and-above and minimum guarantee policies—is so elusive that it has led to unintended consequences in several real-world applications. In some cases, decision makers have completely overlooked this distinction, leaving major loopholes in systems that otherwise meticulously control parameters such as the fraction of reserved units.102102102See the The Hindu op-ed Pathak and Sönmez (2019) for an example of such a loophole when the 103rd Constitutional Amendment Act, 2019 introduced reservations for economically weaker sections of society. Occasionally, politically motivated actors have exploited these loopholes through reforms with significant distributional consequences.103103103See Pathak et al. (2025) for an example of this, when officials of the U.S. Department of Homeland Security reversed the processing sequence of two categories of H-1B visas in response to President Trump’s Buy American and Hire American Executive Order in 2017. In other instances, as illustrated in our next case study, system operators intended to implement either the over-and-above or minimum guarantee policy but ended up implementing the wrong one—either because they were unaware of the distinction or did not know how to operationalize the intended policy, especially in a multi-institution setting.
@@ 9.2 The Demise of Neighborhood Priority in Boston School Choice
Building on Dur et al. (2018b), our next application examines Boston’s neighborhood priority policy and the unintended consequences of its flawed implementation between 2005 and 2013, which ultimately disadvantaged applicants seeking to attend their neighborhood schools.104104104The period from 2005 to 2013 provides a conservative estimate of the duration of the flawed implementation of Boston’s neighborhood priorities. The neighborhood priority policy intended for use under the DA mechanism during this period was originally adopted in 2000 under Boston mechanism. The flawed implementation of this policy, discussed in Section 9.2.2, led to the de facto elimination of neighborhood priorities under the DA mechanism and would have caused similar unintended consequences in the Boston mechanism had an analogous error occurred. However, it remains unclear whether such a flawed implementation of the neighborhood priority policy existed between 2000 and 2005 under the Boston mechanism.
@@ 9.2.3 Phasing Out Neighborhood Priority in Boston: 2013 Reform for Transparency
Though I didn’t know it at the time, one of the regular participants in the public meetings was Peng Shi, an MIT graduate student in operations research who was auditing my market design course at Boston College. Recognizing the gridlock as an opportunity, Shi proposed an entirely new approach during one of the meetings: the “Home-Based” system (Shi, 2015).105105105Later, I learned that Shi had been studying the crisis with an eye toward helping guide a reform as early as Spring 2012. Before the Home-Based system, he also proposed other plans when the EAC elicited suggestions from the wider community in late summer. This system provided each student with a customized list of schools based on their home address, reducing the number of schools they could rank while ensuring that highly-rated options were included.106106106See Shi (2022) for an in-depth exploration of Shi’s broader methodology, which optimizes parameters such as student priority distribution within mechanisms like DA and TTC, parameters that are typically treated as exogenous in design economics. His proposal gained traction, as EAC members were eager to bridge the divide between the committee’s opposing camps.
@@ 10.1 India’s Reservation System and the Quest for Social Justice
Affirmative action for various disadvantaged groups is embedded in the 1950 Constitution of India through articles 14-18.107107107This segment of the Constitution is referred to as the Equality Code in India. In practice, these protective policies are implemented through a reservation system that sets aside a certain fraction of positions in public jobs and seats at public schools for various protected groups. Naturally, the Constitution itself does not provide any rigorous formula to implement these policies. Nevertheless, in a landmark Supreme Court judgment Indra Sawhney vs. Union of India (1992), the justices formulated these provisions in a rigorous way, at least for the more basic versions of the reservation system.108108108Considered the main reference on the reservation system in India, this case is popularly known as the Mandal Commission case.
@@ 10.2 Vertical Reservations, Migration and the Over-and-Above Choice Rule
From a technical point of view, the social identity-based vertical reservation system meant that no individual could be a member of multiple VR-protected groups. While VR protections are also provided to individuals in economic disability since 2019 with the enactment of the 103rd Amendment,109109109This more recent vertical category is called Economically Weak Segments (EWS). members of the earlier and social identity-based VR-protected groups are excluded from its scope,110110110This exclusion is highly controversial in India. See Sönmez and Ünver (2022) for a case study in minimalist market design that analyzes this exclusion in the 103rd Amendment. thereby still maintaining the non-overlapping structure of VR-protected groups in the country.
@@
A subtle but important aspect of the above mandate is its reliance on a basic technical tool commonly referred to as migration in Indian legal terminology.111111111The concept of migration is also known as mobility in Indian legal terminology.
@@ 10.3 Horizontal Reservations, Adjustments and the Minimum Guarantee Choice Rule
This passage leaves unclear what exactly the “necessary adjustments” refer to and how they are to be made. The ambiguity surrounding these adjustments and the implementation of the HR policy was later clarified by the Supreme Court in Anil Kumar Gupta vs. State of U.P. (1995).112112112The formulation of the HR policy in Anil Kumar Gupta vs. State of U.P. (1995) is further clarified in a subsequent Supreme Court judgment, Rajesh Kumar Daria vs. Rajasthan Public Service Commission (2007). In this judgment, the Supreme Court further elaborated two key characteristics of the HR policy that distinguish it from the VR policy:
@@
Possibly to avoid complications arising from overlaps between “property rights” gained through these two types of affirmative action policies, the justices also recommended implementing the HR policy within each vertical category, including the open category.113113113This mode of HR policy is referred to as compartmentalized horizontal reservations in Anil Kumar Gupta vs. State of U.P. (1995). For example, if there are horizontal reservations for women, they are provided with preferential access for a fraction of the positions in each category (i.e., open, SC, ST, OBC, etc.) separately.
@@
Their procedure begins by disregarding the HR policy in the first two steps and focusing solely on the VR policy. In the first step, open positions are tentatively allocated using SSD, followed by the allocation of VR-protected positions to the remaining eligible individuals, again using SSD, in the second step.114114114Note that these two steps together form the over-and-above choice rule.
@@
Except for the recent EWS category, a member of a VR-protected category is not required to declare membership in their category.115115115This contrasts with the documentation needed to prove ineligibility for social identity (i.e., caste or tribal) based VR-protected categories, which is a prerequisite for claiming membership in the EWS category. This typically involves documentation of forward caste membership in India. By claiming their VR protections, an individual—unless a meritorious reserved candidate—loses their open category HR protections under the SCI-AKG choice rule. Thus, this mechanism violates a form of the incentive compatibility axiom, first formulated by Aygün and Bó (2021) in the context of affirmative action at Brazilian colleges.
@@ 10.5 External Validity for Minimalist Market Design: Life Imitates Science in a Landmark Judgment
Before its publication, the initial draft of Sönmez and Yenmez (2022) was circulated in March 2019 as a Boston College working paper Sönmez and Yenmez (2019). In this work, we established that the Supreme Court-mandated SCI-AKG choice rule had been consistently undermining intended affirmative action measures in real-world practices by failing to satisfy the no justified envy axiom.116116116In Sönmez and Yenmez (2019), the SCI-AKG choice rule is referred to as the SCI-VHR choice rule, and the no justified envy axiom is referred to as elimination of justified envy. We also documented how this failure sparked numerous litigations between 1995 and 2020, causing recurring disruptions in the country’s recruitment processes.
@@
Motivated by this case and several others that contradicted the ruling in Anil Kumar Gupta vs. State of U.P. (1995), we proposed the 2SMG choice rule in Sönmez and Yenmez (2019) as a minimalist reform of the SCI-AKG choice rule.117117117As documented in Sönmez and Yenmez (2019, 2022), the challenges arising from the failures of the Supreme Court-mandated SCI-AKG choice rule extended far beyond conflicting with high court judgments that upheld choice rules satisfying the no justified envy axiom. In many instances, state institutions using the SCI-AKG choice rule were challenged in court by VR-protected candidates due to cases of justified envy. Some of these cases were ruled in favor of the state, while others favored the plaintiffs. Additionally, there were situations where VR-protected candidates were denied their open-category HR protections even though they did not claim their VR protections. Court decisions on these challenges also varied, sometimes favoring the candidates and other times not. Moreover, in other cases, VR-protected candidates were denied open-category HR-protected positions even when no positions were reserved for their VR-protected categories.
@@
    Awarding $i$ an open-category position instead, with or without replacing another individual, does not increase the number of HR-protected positions honored within the open category.118118118Prior to Saurav Yadav vs. State of U.P. (2020), unlike the first two conditions, the necessity of this third condition as a prerequisite for the assignment of VR-protected positions remained unclear.
@@
Therefore, while the justices have not explicitly mandated the 2SMG choice rule in Saurav Yadav vs. State of U.P. (2020), they have, in effect, mandated it in the country, as no other choice rule satisfies the explicit mandates of the judgment.119119119The 2SMG choice rule is mandated in the state of Gujarat under the High Court judgment Tamannaben Ashokbhai Desai vs. Shital Amrutlal Nishar (2020).
Aligned with the version of the problem reviewed in Saurav Yadav vs. State of U.P. (2020), Theorem 12 pertains to the basic version of the problem, where all positions are identical and no individual belongs to more than one HR-protected group.120120120Strictly speaking, Saurav Yadav vs. State of U.P. (2020) does not rule out the possibility of individuals belonging to multiple HR-protected groups; it simply does not address this scenario.
@@
10.6 Overlapping Horizontal Reservations
While Persons with Disabilities (PwD) is the only HR-protected group recognized nationwide in India by a Supreme Court judgment,121121121HR protections for Persons with Disabilities were established in the Supreme Court judgment Union of India vs. National Federation of the Blind (2013). other groups, such as Women (W), receive HR protections in several states through mandates issued by High Courts or state governments. As a result, unlike VR-protected groups, individuals can often belong to multiple HR-protected groups in practical applications.